What are the odds against drawing a red ten, in a standard 52 card deck?
Answer by stanbon(75887)

What are the odds against drawing a red ten, in a standard 52 card deck?
-------------
P(red ten) = 2/52
----
P(not a red ten) = 50/52
-----
Odds against drawing red ten = P(not a red ten)/P(red ten) = 50:2 = 25:1
------------------------
Cheers,
Stan H.
================